The impedance in a circuit is the measure of how much a circuit impedes (hinders) the flow of current through it. The impedance of a resistor is Z[1]=3+15i ohms and the impedance of another resistor is Z[2]=4+6i ohms. When the twp resistors are connected in parallel, the total impedance is given by
Z[r]= Z[1]·Z[2]/Z[1]+Z[2]

Find the total impedance

Answer by stanbon(75887)

The impedance in a circuit is the measure of how much a circuit impedes (hinders) the flow of current through it. The impedance of a resistor is Z[1]=3+15i ohms and the impedance of another resistor is Z[2]=4+6i ohms. When the twp resistors are connected in parallel, the total impedance is given by
Z[r]= {Z[1]·Z[2]}/{Z[1]+Z[2]}
---
z(r) = [4+6i+3+15i]/[(4+6i)(3+15i)]
------
= [7+21i]/[(12-90)+(18+60)i]
------
= (7+21i)/(-78+78i)
--------------------------
Cheers,
Stan H.
====================